2\5x - 1\2y -5=0 and 0.07x + 0.06y-200=0
Answer by mananth(16949)

2\5x - 1\2y -5=0
4x-5y=50.........................1
0.07x + 0.06y-200=0
multiply by 100
7x+6y=20000......................2
4 x + -5 y = 50 .............1
7 x + 6 y = 20000 .............2

multiply (1)by 6
Multiply (2) by 5
24 x + -30 y = 300
35 x + 30 y = 100000

Add the two equations

59x= 100300

/59

x=1700

plug value of x in (1)

4 x + -5 y = 50
6800 + -5 y = 50
-5 y = 50 + -6800
-5 y = -6750
y = 1350